Crane, Inc., is considering investing in a new production line for eye drops. Other than investing in the equipment, the company needs to increase its cash and cash equivalents by $10,000, increase the level of inventory by $28,000, increase accounts receivable by $25,000, and increase accounts payable by $5,000 at the beginning of the project. Crane will recover these changes in working capital at the end of the project 12 years later. Assume the appropriate discount rate is 11 percent. What are the present values of the relevant investment cash flows? (Do not round intermediate calculations. Round answer to 2 decimal places, e.g. 15.25.)
Answer #1

Net Working Capital = Cash + Inventory + AR - AP = 10,000 + 28,000 + 25,000 - 5,000 = 58,000

NWC will be required to invest in year 0 and the same will be redeemed in year 12

=> Present Value (PV) = NWC + NWC / (1 + r)^n = -58,000 + 58,000 / (1 + 11%)^12 = -41,421.20